OpenStudy (ria23):

How would I graph x - 2y >= 10?

OpenStudy (ria23):

\[x - 2y \ge 10\]

zepdrix (zepdrix):

Yhu again -_-

OpenStudy (ria23):

Haha, ^u^ Hi!

zepdrix (zepdrix):

Ok so like, you remember how to graph stuff when it's \(\Large\rm y=(stuff)\). So first try to get it into that form.

OpenStudy (ria23):

XD I got that much... Not really sure how to tho...

zepdrix (zepdrix):

\[\Large\rm x - 2y \ge 10\]Mmm I guess we'll start by subtracting x from each side,\[\Large\rm -2y \ge -x+10\]Mmmmm what next?

OpenStudy (ria23):

Um... Yhu have to get y by itself,,, So yhu divide both sides by negative 2 right?

zepdrix (zepdrix):

ye. Do you remember what happens when you `divide a negative across an inequality`?

OpenStudy (ria23):

Does is... Switch signs?

OpenStudy (ria23):

it*

zepdrix (zepdrix):

Yeh :U It switches directions.

zepdrix (zepdrix):

\[\Large\rm -2y \ge -x+10\]Becomes:\[\Large\rm y \le \frac{1}{2}x-5\]

zepdrix (zepdrix):

x/2 is the same as (1/2)x. I wrote it that way so we can identify the slope easier.

zepdrix (zepdrix):

So first think about \(\Large\rm y=\dfrac{1}{2}x-5\). Understand how to graph that? Do it -_-

OpenStudy (ria23):

xD So bossy.

zepdrix (zepdrix):

Mmm I wish you wouldn't draw your axes so weirdddd :( We can't see most of the line... grr

OpenStudy (ria23):

XD sorry... I just kind of drew it and added as I went

zepdrix (zepdrix):

Mmmm ok how bout we use this -_-